# Stars and quark stars in bumblebee gravity

### v2

@@ IV Final remarks
Recently, three works on bumblebee gravity and stars drawn our attention. The first one Ji:2024aeg builds an interior solution from another exterior solution in bumblebee gravity, namely the spacetime obtained in Bertolami and Páramos Bertolami . In our case, as mentioned before, the exterior solution is the Schwarzschild-like metric of Casana et al. Casana . In this case, there is no overlapping results. Regarding the two other papers, those authors tried to obtain the interior solution of a star from the Schwarzschild-like metric. The modified TOV is incorrect in either cases. In Ref. Liu , the polytropic fluid studied is not solution of all modified field equations, and in Ref. Panotopoulos:2024jtn even the exterior solution is not solution of the field equations in the bumblebee context.
@@ Acknowledgments
JCSN thanks Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq), Brazil, (Grant No. 170579/2023-9) for the financial support and the ICT-Unifal for the kind hospitality. F.G.G. is supported by CNPq through 306762/2021-8.

### v3

@@ V Final remarks
Another interesting point regards the stability analysis. The $M-\rho_{c}$ curve—from the $\frac{dM}{d\rho_{c}}>0$ stability criterion—indicates stability for the bumblebee stars with quark matter. Moreover, large values of the Lorentz-violating parameter $\ell$ provide larger mass-radius relations, but the range of allowed and stable configurations decreases with that parameter.
@@ Acknowledgments
JCSN thanks Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq), Brazil, (Grant No. 170579/2023-9) for the financial support and the ICT-Unifal for the kind hospitality. F.G.G. is supported by CNPq through 306762/2021-8. The authors also thank two anonymous referees for valuable comments.